OpenStudy (anonymous):

Stephanie has 23 coins in her piggy bank. She only puts dimes and pennies inside. If she has $1.58, how many dimes does she have?

OpenStudy (anonymous):

Set this up as a system of equations where x = the number of dimes and y = the number of pennies If she has a total of 23 coins then one equation is \[x+y=23\] And if the total is $1.58 the other equation is \[10x + y = 158\] subtract the two equations and solve for x \[(10x - x) + (y - y) = (158-23)\] \[9x = 135\] \[x = 15\] Thus, Stephanie has 15 dimes and 8 pennies.